OpenStudy (anonymous):

Modular arithmetic. 21^20 = 1 (mod 100) From here it's simple to get that 21^(20+1) = 21 (mod 100) But what about 21^(20-1) = ? (mod 100) Is there a simple method for this?

Parth (parthkohli):

\[21^{15} \cdot 21^4 \equiv 1 \cdot 81 \equiv 81 \pmod{100} \]

OpenStudy (anonymous):

Thanks, but I'm looking for something more general. Like: \[p^ {n+1} = p \mod 100\] Where p^n = 1 mod 100

OpenStudy (anonymous):

\[21==21\]

Parth (parthkohli):

You have no way but to use the modulos.

OpenStudy (anonymous):

\[21^2==41\]

OpenStudy (anonymous):

\[21^3==61\]

Parth (parthkohli):

\[21^4 \equiv 81\]

Parth (parthkohli):

Yes, that is what I used to determine all the mods: pattern recognition. :-)

OpenStudy (anonymous):

no until you reach 1 later on you can get the powr of 1

Parth (parthkohli):

And write it in some crappy notation so that people think you are smart.

Parth (parthkohli):

Repeats every four terms.

OpenStudy (anonymous):

\[21^5=1\]

OpenStudy (anonymous):

Well what if I have something a bigger repetition pattern? I don't want to brute force the problem.

Parth (parthkohli):

You just have to find a power which is \(1\) or \(-1\) that thing. (thanks @terenzreignz!)

Parth (parthkohli):

\(1\) or \(-1\) modulo*

Parth (parthkohli):

You always have Fermat's Little Theorem!

terenzreignz (terenzreignz):

Cute :) Then again, you could also express it as (20 + 1)^20 (mod 100) If that's any easier :D

Parth (parthkohli):

\[a^{b - 1} \equiv 1 \pmod{b}\]